OpenStudy (anonymous):

find f(g(x)) and g(f(x)) and determine wether the pair of functions f and g are inverse of each other. f(x)=8x+8 and g(x)=x-8/8 f(g(x))=

myininaya (myininaya):

\[f(x)=8x+8, g(x)=\frac{x-8}{8}\] \[f(g(x))=f(\frac{x-8}{8})=8(\frac{x-8}{8})+8=x-8+8=x+0=x\] \[g(f(x))=g(8x+8)=\frac{(8x+8)-8}{8}=\frac{8x}{8}=x\] f and g are inverses of each other

OpenStudy (anonymous):

You do the "inside first" so you the value of g(x) into the function (f(x) And vice versa for the other one. Do you understand?

OpenStudy (anonymous):

you can also try it with numbers to get an idea of what is going on \[f(2)=8\times 2+8=16+8=24\] \[g(24)=\frac{24-8}{8}=\frac{16}{8}=2\]

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

so if f sends 2 to 24, g sends 24 to 2 and do on

myininaya (myininaya):

x is a number lol

OpenStudy (anonymous):

well you could try it with \[\pi\]

myininaya (myininaya):

or \[e^\pi\]

OpenStudy (anonymous):

show off

OpenStudy (anonymous):

btw you know from elementary algebra that if you want to solve \[8x+8=24\] for x, you would do the following steps: a) subtract 8 b) divide by 2 that is what the inverse function says to do